Urea (H2NCONH2) is used extensively as a nitrogen source in fertilizers. It is produced commercially from the reaction of ammonia and carbon dioxide:

2NH3 (g) + CO2 (g)  H2NCONH2 (s) + H2O (g)
Ammonia gas at 223oC and 90 atm flows into a reactor at a rate of 500 L/min. Carbon dioxide at 223oC and 45 atm flows into the reactor at a rate of 600 L/min. What mass of urea is produced per minute by this reaction assuming 100% yield?
1
Expert's answer
2020-03-18T12:54:29-0400

Applying "pV=nRT"

"n=\frac{pV}{RT}"


"n=\frac{90\times500}{496\times0.082} =1106 mole/min" for ammonia


"n=\frac{45\times600}{496\times0.082} =664mole /min" for carbon-dioxide

Here if we compare the stochiometric coefficients of ammonia and carbon dioxide then it is 2:1

then here the limiting reagent is ammonia

As the ratio of stochiometric coefficients of ammonia and urea is in the ratio 2:1

So urea produced will be half of ammonia=1106/2 mole/min =553 mole/min

urea produced in mass = 553*60 =33180grams/min=33.18 kg/min